OpenStudy (anonymous):

A spherical balloon with radius r inches has volume V(r) = 4 3 πr3. Find an expression that represents the amount of air required to inflate the balloon from a radius of r inches to a radius of r + 3 inches. (Express your answer in terms of π and r.)

OpenStudy (goformit100):

First of all @bnblack : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

you want to find the change in volume: V = V(r+3) - V(r)

OpenStudy (anonymous):

Okay thank you very much!

OpenStudy (anonymous):

V(r) is already given. To find V(r+3) plug (r+3) instead of r in V(r) \[V(r+3)=\frac{ 4 }{ 3 }π(r+3)^3\] let me know if you need help expanding (r+3)^3

OpenStudy (anonymous):

glad i could help :)